The length of a rectangular garden is 3 meters more than its width. Find the dimensions of the garden if its area is 18 square meters
Answer by mananth(16949)

width =x
length = x+3
Area = l*w
Area =x(x+3) =18
x^2+3x-18=0
x^2+6x-3x-18=0
x(x+6)-3(x+6)=0
(x-3)(x+6)=0
so x=3m width
length = 3+3=6m